If you wish to accumulate $249,000 in 6 years, how much must you deposit today in an account that pays a quoted annual interest rate of 13% with quarterly compounding of interest?

$119,599.31
$115,568.08
$159,067.09
$105,282.52
$198,534.86
Answer #1

Quoted Interest Rate = 13 % per annum or (13/4) = 3.25 % per quarter

Targeted Future Value = $ 249000 and Tenure = 6 years or (6 x 4) = 24 quarters

Let the required current deposit be $ P

P x (1.0325)^(24) = $ 249000

P = 249000 / (1.0325)^(24) = $ 115568.08

Hence, the correct option is (b)